OpenStudy (anonymous):

How do I solve ln(8 - x) = 0

hartnn (hartnn):

ok, as \(\ln a=b \implies a=e^b \) here, you get 8-x = e^0 = 0 so can you solve 8-x=1 now ?

OpenStudy (anonymous):

thx :)

hartnn (hartnn):

welcome ^_^